OpenStudy (anonymous):

A ball is launched from a height of 6 feet with an initial velocity of 92 feet per second. What is the maximum height the ball will reach? When will the ball hit the ground?

OpenStudy (anonymous):

So would it be 738,881 is the. Adium height?

OpenStudy (anonymous):

*maximum

OpenStudy (aum):

\(y = ut - \frac 12 gt^2 + h = 92t - \frac 12 * 32 * t^2 + 6\) \(y = -16t^2 + 92t + 6\) -------- (1) The graph of y vs t will be a parabola that opens downward. The y-coordinate of the vertex gives the maximum height. The x-coordinate of the vertex of the general equation y = ax^2 + bx + c is -b/(2a). Here, b = 92 and a = -16. b/(2a) = -92 / (2 * -16) = 92/32 = 23/8 = 2.875 seconds. Put t = 2.875 in (1) and calculate y and that will be the maximum height. Find the roots of the equation and that will give the time it takes for the bill to hit the ground. You can also find the total time by doubling the time it takes to reach the maximum height which we calculated earlier as 2.875 s. So double that.

OpenStudy (aum):

No.

OpenStudy (aum):

put t = 2.875 in \(y = -16t^2 + 92t + 6 \) and calculate y.

OpenStudy (anonymous):

264.2 ft

OpenStudy (aum):

Are you just throwing out numbers from the choices without actually doing the calculations?

OpenStudy (anonymous):

No I am doing it I just don't get it